OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 8, x = 2. Find y when x = 8.

OpenStudy (shubhamsrg):

well y= k/2x // where k is constant so y1 = k/2x1 and y2 = k/2x2 divide both eqns and substitute values//